Find the center of the mass of four 2 kg bodies placed at each corner of a square of side 1.5 m. Show the given, required, equations, solutions and answer.

1
Expert's answer
2021-11-04T16:20:19-0400


Let the left bottom body be at the origin.



"4mX=m(0+0+1.5+1.5)\\X=0.75\ m
\\4mY=m(0+0+1.5+1.5)\\Y=0.75\ m\\"

COM is at center of square.